OpenStudy (taiofchi):

Jebb is the tallest player on the basketball team. He is 1 1/2 times as tall as the shortest girl in the sixth grade, who is 4 1/4 feet tall. How tall is Jebb? Please help me.

OpenStudy (ryuga):

1 1/2 times 4 1/4 change to improper fractions (3/2)*(17/4) (3*17)/(2*4)=51/8=6 3/8 feet tall 3/8 of 12 inches is (3/8)*12=36/8=4 4/8=4.5 inches Jebb is 6 feet 4.5 inches tall

OpenStudy (taiofchi):

Okay...... Thanks ;D

OpenStudy (ryuga):

Anymore?

OpenStudy (taiofchi):

Nope

OpenStudy (taiofchi):

thanks again that was a question i got stuck and confused on...